THE DEGREE OF THE SECANT VARIETY AND THE JOIN OF MONOMIAL CURVES

KRISTIAN RANESTAD

Abstract. A monomial curve is a curve parametrized by monomials. The degree of the secant variety of a monomial curve is given in terms of the sequence of exponents of the monomials defining the curve. Likewise, the degree of the join of two monomial curves is given in terms of the two sequences of exponents.